OpenStudy (anonymous):

d. find the derivative how to set up y=log4(3x–2)3(7x2+5) is the answer 1/[9(3x-2)^2(14x)]ln4

OpenStudy (anonymous):

log base 4

OpenStudy (anonymous):

Close, but you need to use the chain rule and the product rule.

OpenStudy (anonymous):

On the inside of the log

OpenStudy (anonymous):

which part is the chain rule for

OpenStudy (anonymous):

I'll get ya started: \[\frac{d}{dx}log_4(f(x)) = \frac{1}{f(x)ln4}\cdot \frac{d}{dx}f(x)\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

So in this case f(x) is?

OpenStudy (anonymous):

Oh wait a min

OpenStudy (anonymous):

You need to be more clear with your brackets. Is the question this: \[log_4[(3x–2)^3(7x2+5)]\]or\[log_4([3x–2]^3)\cdot (7x2+5)\]or\[log_4([3x–2])^3\cdot (7x2+5)\]

OpenStudy (anonymous):

dont tell me i was right

OpenStudy (anonymous):

yes first one

OpenStudy (anonymous):

No you weren't right for sure. I'm just not sure what you're trying to solve. But either way you'd need the product rule and you don't have it.

OpenStudy (anonymous):

So it's the log base 4 of the whole thing.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Actually then the easiest thing to do is break up the product into a sum of logs.

OpenStudy (anonymous):

\[log_4((a)^3(b)) = log_4(a^3) + log_4(b)\]

OpenStudy (anonymous):

log4[(3x–2)3(7x2+5)] just for reference page is running out

OpenStudy (anonymous):

Then differentiate term by term

OpenStudy (anonymous):

using the chain rule. Also you can move that exponent inside into a multiplicative constant outside: \[log_b(a^k) = k(log_b(a))\]

OpenStudy (anonymous):

these are just standard logarithm rules, but they help make things easier so you don't have to fiddle as much with chain and product rules. You get the same result, but much faster.

OpenStudy (anonymous):

the last part is the answer